Show that if S and T are linear

transformations on a finite dimensional

vector space, then rank (ST)<= rank (S).


1
Expert's answer
2021-07-01T11:51:45-0400

As "V" is a finite dimensional vector space, let dimension of "V \ is \ n" .

According to question,

Let , "T:V\rightarrow V" and "S:V\rightarrow V" are two linear transformation from a finite dimensional vector space "V" to itself.

Then "S\circ T :V \rightarrow V" is a linear transformation.

Now, "dim(ImT)=rank(T) \ ,\ dim(ImS)=rank(S)"

and "dim(Im(ST))=dim(Im(S\circ T))=rank (ST)."

"dim(Ker (S))=nullity (S),dim(ker(T))=nullity(T)"

and "dim(ker(ST))=nullity (ST)" .

Note: Generally,"\ S\circ T" written as "ST" .

Now ,"Ker(ST)=\{x\in V :(ST)(x)=0 \}" .

Where is a zero vector of "V."

"=\{ x\in V:S(T(x))=0 \}"

Now , if "x\in ker(T)" then

"T(x)=0 \ \implies S(T(x))=S(0)=0" .

Hence, "x\in Ker(ST)."

Therefore,"Ker(T)\sube Ker(ST)."

Thus,"dim(Ker(T))\leq dim(Ker(ST))"

I,e,"nullity (T)\leq nullity (ST)...............(1)"

Now , from the rank-nullity theorem ,

"rank(T)+nullity(T)=rank(ST)+nullity (ST)=n"

Now , from (1) we concluded ,

"rank(ST)\leq rank(S)" .